. An investor expects to receive Rs 572750, at the end of 20 years to keep his retirement life happy. Discuss how much money he should invest in lump sum to meet his expectation at the given prevailing interest rate of 10% per annum. Discuss the concept of time value of money you applied here.


3.b. How much wealth Rs 20000 would produce if it’s invested in a fixed deposit for a tenure of 12 years at an interest of 8% per annum. Also, discuss the concept of compounding and

calculating the future value of money.



1
Expert's answer
2021-10-04T10:18:09-0400

a)Time value of money concept state that value of money in future would not be same that in present. The money will discount at over the time.

In order calculate value of money today, we need to use the present value formula

Present value =Future Value/(1+interest rate)number of years 

Future value =572750

Interest rate =10%

Number of years =20

"Present \space Value =572750/(1+0.10)^{20}\\

 =\frac{572750}{6.7275}\\

 =85135.64"

Hence money need to invest today is 85135.64

We applied here the present value concept here in which future money is discounted at appropriate rat


b)

A=final amount

P=initial principal balance

i=interest rate

nt=number of times interest applied per time period

"A=P(1+\frac{r}{n})^{nt}\\A=20000(1+\frac{0.008}{12})^{12}\\=\$ 20160.59"

Compounding concept determine what will we get tomorrow if we invest a certain sum. Future value of money tables the concept